# Convergence Rates for the Maximum A Posteriori Estimator in PDE-Regression Models with Random Design

### v1

@@ 1. C’7)
    Inverse Continuity Modulus: Let $f_{1},f_{2}\in\mathcal{F}$ for some $\alpha>2+\frac{d}{2}$. Then there exists a constant $\operatorname{B}\in\mathbb{R}_{>0}$, such that $\left\|f_{1}\right\|_{C^{1}(\mathcal{O})}\lor\left\|G^{D}(f_{2})\right\|_{C^{2}(\mathcal{O})}\leq\operatorname{B}$. Lemma 24 in Nickl et al. [2020] provides the existence of a finite constant $c_{1}=c_{1}(g_{\min},f_{\min},\operatorname{B},\mathcal{O},d)\in\mathbb{R}_{>0}$, such that
    $\left\|f_{1}-f_{2}\right\|_{\mathbb{L}^{2}(\mathcal{O})}\leq c_{1}\cdot\left\|f_{2}\right\|_{C^{1}(\mathcal{O})}\cdot\left\|G^{D}(f_{1})-G^{D}(f_{2})\right\|_{H^{2}(\mathcal{O})}.$
    By an application of Eq. A.9, we have further
    $\left\|G^{D}(f_{1})-G^{D}(f_{2})\right\|_{H^{2}(\mathcal{O})}\leq c_{2}(\alpha,\mathcal{O})\cdot\left\|G^{D}(f_{1})-G^{D}(f_{2})\right\|_{H^{\alpha+1}(\mathcal{O})}^{\frac{2}{\alpha+1}}\cdot\left\|G^{D}(f_{1})-G^{D}(f_{2})\right\|_{\mathbb{L}^{2}(\mathcal{O})}^{\frac{\alpha-1}{\alpha+1}}.$
    Moreover, by an application of Lemma 23 in Nickl et al. [2020], we have
    $\left\|G^{D}(f_{1})-G^{D}(f_{2})\right\|_{H^{\alpha+1}(\mathcal{O})}^{\frac{2}{\alpha+1}}\leq c_{3}\cdot(1+\left\|f_{1}\right\|_{H^{\alpha}(\mathcal{O})}^{2\alpha}\lor\left\|f_{2}\right\|_{H^{\alpha}(\mathcal{O})}^{2\alpha})$
    with a finite constant $c_{3}=c_{3}(\alpha,d,\mathcal{O},f_{\min},\left\|g\right\|_{\alpha-1})\in\mathbb{R}_{>0}$. Thus, we have
    $\left\|f_{1}-f_{2}\right\|_{\mathbb{L}^{2}(\mathcal{O})}\leq c_{4}\cdot(1+\left\|f_{1}\right\|_{H^{\alpha}(\mathcal{O})}^{2\alpha}\lor\left\|f_{2}\right\|_{H^{\alpha}(\mathcal{O})}^{2\alpha})\cdot\left\|G^{D}(f_{1})-G^{D}(f_{2})\right\|_{\mathbb{L}^{2}(\mathcal{O})}^{\frac{\alpha-1}{\alpha+1}}$
    for a constant $c_{4}=c_{4}(\alpha,d,\mathcal{O},f_{\min},g_{\min},\operatorname{B},\left\|g\right\|_{\alpha-1})\in\mathbb{R}_{>0}$. Now, let $\operatorname{M}\in\mathbb{R}_{>0}$ be arbitrary and assume that $\left\|f_{1}\right\|_{H^{\alpha}(\mathcal{O})}+\left\|f_{2}\right\|_{H^{\alpha}(\mathcal{O})}\leq\operatorname{M}$, such that $\left\|G^{D}(f_{1})-G^{D}(f_{2})\right\|_{\mathbb{L}^{2}(\mathcal{O})}\leq\delta$ for $\delta\in\mathbb{R}_{>0}$ small enough. By observing that $\left\|f_{1}\right\|_{C^{1}(\mathcal{O})}\lor\left\|G^{D}(f_{2})\right\|_{C^{2}(\mathcal{O})}\leq c(\operatorname{M},\alpha,\left\|g\right\|_{H^{\alpha-1}(\mathcal{O})})$, we obtain
    $\left\|f_{1}-f_{2}\right\|_{\mathbb{L}^{2}(\mathcal{O})}\leq c_{5}\cdot\delta^{\frac{\alpha-1}{\alpha+1}}$
    with $c_{5}=c_{5}(\alpha,d,\mathcal{O},f_{\min},g_{\min},\operatorname{M},\left\|g\right\|_{\alpha-1})\in\mathbb{R}_{>0}$. This shows the claim with $\tau=\frac{\alpha-1}{\alpha+1}\in(0,1)$.
Proof of 9 (Theorem 3.2).
@@
    Thus, for every $\bar{c}\in\mathbb{R}_{>0}$ there exists a constant $c_{1}=c_{1}(\alpha,f_{\min},\left\|g\right\|_{\infty},d,\mathcal{O},\sigma)\in\mathbb{R}_{>0}$ sufficiently large, such that
    $\mathbb{P}_{f_{\mathfrak{o}}}^{N}\left(\mathfrak{D}_{\mathfrak{r}}(\hat{f}_{\textsc{MAP}},f_{\mathfrak{o}})\geq c_{1}(\mathfrak{r}_{N}^{2}\left\|\Psi^{-1}\circ f_{\mathfrak{o}}\right\|_{H^{\alpha}(\mathcal{O})}^{2}+m^{2})\right)\leq c_{2}\exp\leq\left(-\bar{c}m^{2}N\right)$
    for all $m\geq\mathfrak{r}_{N}$ satisfying Eq. C4 and Eq. C5. We identify $f_{\mathfrak{o}}\in\mathcal{F}$ with some $\theta_{\mathfrak{o}}=\Psi^{-1}\circ f_{\mathfrak{o}}\in H^{\alpha}_{c}(\mathcal{O})$ and conclude $\left\|\Psi^{-1}\circ f_{\mathfrak{o}}\right\|_{H^{\alpha}(\mathcal{O})}\leq c_{3}(\Psi,\operatorname{L})\in\mathbb{R}_{>0}$ uniformly over all $f_{\mathfrak{o}}\in\mathcal{F}(\operatorname{L})$. Choosing $\operatorname{M}\geq 1\lor c_{3}$ and $m:=\operatorname{M}\mathfrak{r}_{N}\geq\mathfrak{r}_{N}$, such that $c_{1}(\mathfrak{r}_{N}^{2}\left\|\Psi^{-1}\circ f_{\mathfrak{o}}\right\|_{H^{\alpha}(\mathcal{O})}^{2}+m^{2})\leq 2c_{1}\operatorname{M}^{2}\mathfrak{r}_{N}^{2}$. Thus,
    $\mathbb{P}_{f_{\mathfrak{o}}}^{N}\left(\mathfrak{D}_{\mathfrak{r}}(\hat{f}_{\textsc{MAP}},f_{\mathfrak{o}})\geq 2c_{1}\operatorname{M}^{2}\mathfrak{r}_{N}^{2}\right)\leq c_{2}\exp\leq\left(-\bar{c}M^{2}N\mathfrak{r}_{N}^{2}\right).$
@@
Proof of 10 (Theorem 3.3).
Observing that we have shown the analogue results to the random design regression model, the proof of this theorem follows exactly along the lines of corresponding proofs in Nickl et al. [2020]. Hence, we are here only providing a scetch of the proof and ommit the details, which are let for the interested reader.
  1. i)
    By an applycatio of Eq. A.9, we obtain
    $\left\|G^{D}(\hat{f}_{\textsc{MAP}})-G^{D}(f_{\mathfrak{o}})\right\|_{H^{\beta}(\mathcal{O})}\leq c_{1}\cdot\left\|G^{D}(\hat{f}_{\textsc{MAP}})-G^{D}(f_{\mathfrak{o}})\right\|_{\mathbb{L}^{2}(\mathcal{O})}^{\frac{\alpha+1-\beta}{\alpha+1}}\cdot\left\|G^{D}(\hat{f}_{\textsc{MAP}})-G^{D}(f_{\mathfrak{o}})\right\|_{H^{\alpha+1}(\mathcal{O})}^{\frac{\beta}{\alpha+1}}.$
    While the second factor can be upper bounded by regularity estimates for $G^{D}$, the first factor is already indicating the correct exponent. Indeed, the first factor can be upper bounded by $\mathfrak{D}_{\mathfrak{r}}(\hat{\theta}_{\textsc{MAP}},\theta_{\mathfrak{o}})$. Decomposing the underlying probability space into slices
    $A_{0}:=\left\{\mathfrak{D}_{\mathfrak{r}}(\hat{f}_{\textsc{MAP}},f_{\mathfrak{o}})<\delta_{N}\right\}\text{ and }A_{j}:=\left\{\mathfrak{D}_{\mathfrak{r}}(\hat{f}_{\textsc{MAP}},f_{\mathfrak{o}})\in(2^{j-1}\delta_{N},2^{j}\delta_{N}]\right\}$
    for some $\delta_{N}\approx\mathfrak{r}_{N}$, the corresponding expectation can be upper bounded by
    $\mathbb{E}_{f_{\mathfrak{o}}}^{N}\left[\left\|G^{D}(\hat{f}_{\textsc{MAP}})-G^{D}(f_{\mathfrak{o}})\right\|_{H^{\beta}(\mathcal{O})}\right]\lesssim\sum_{j\in\mathbb{N}_{0}}\mathbb{E}_{f_{\mathfrak{o}}}^{N}\left[\mathds{1}_{A_{j}}\mathfrak{D}_{\mathfrak{r}}(\hat{\theta}_{\textsc{MAP}},\theta_{\mathfrak{o}})^{\frac{\alpha+1-\beta}{\alpha+1}}\left(1+\mathfrak{D}_{\mathfrak{r}}(\hat{\theta}_{\textsc{MAP}},\theta_{\mathfrak{o}})^{\alpha^{2}\beta}\right)\right].$
    Using the upper bound of $\mathfrak{D}_{\mathfrak{r}}(\hat{\theta}_{\textsc{MAP}},\theta_{\mathfrak{o}})$ on each event $A_{j}$, as well as the concentration inequality of Theorem 3.2, we obtain
    $\mathbb{E}_{f_{\mathfrak{o}}}^{N}\left[\left\|G^{D}(\hat{f}_{\textsc{MAP}})-G^{D}(f_{\mathfrak{o}})\right\|_{H^{\beta}(\mathcal{O})}\right]\lesssim\delta_{N}^{\frac{\alpha+1-\beta}{\alpha+1}}\cdot(1+o(N^{-\frac{1}{2}})),$
@@
  2. i)
    Similiarly, as in the proof of C’7) before, we can upper bound
    $\left\|\hat{f}_{\textsc{MAP}}-f_{\mathfrak{o}}\right\|_{\mathbb{L}^{2}(\mathcal{O})}\lesssim\left\|\hat{f}_{\textsc{MAP}}\right\|_{C^{1}(\mathcal{O})}\cdot\left\|G^{D}(\hat{f}_{\textsc{MAP}})-G^{D}(f_{\mathfrak{o}})\right\|_{H^{2}(\mathcal{O})}.$
    Thus, by an application of Eq. A.9 with $\beta=2$, we obtain
    $\left\|G^{D}(\hat{f}_{\textsc{MAP}}-G^{D}(f_{\mathfrak{o}})\right\|_{H^{2}(\mathcal{O})}\lesssim\left\|G^{D}(\hat{f}_{\textsc{MAP}}-G^{D}(f_{\mathfrak{o}})\right\|_{H^{1+\alpha}(\mathcal{O})}^{\frac{2}{1+\alpha}}\cdot\left\|G^{D}(\hat{f}_{\textsc{MAP}})-G^{D}(f_{\mathfrak{o}})\right\|_{\mathbb{L}^{2}(\mathcal{O})}^{\frac{\alpha-1}{\alpha+1}}.$
    The proof now follows the same idea as for the prediction error, to obtain an upper bound of the form
    $\mathbb{E}_{f_{\mathfrak{o}}}^{N}\left[\left\|\hat{f}_{\textsc{MAP}}-f_{\mathfrak{o}}\right\|_{\mathbb{L}^{2}(\mathcal{O})}\right]\lesssim\delta_{N}^{\frac{\alpha-1}{\alpha+1}}\cdot(1+o(N^{-\frac{1}{2}})),$
@@
   Auxiliary results
The following lemma was first proven in Nickl and Wang [2024] (see Lemma 3.12) and is a key tool in the proof of Theorem 2.6.
Lemma § A.7 (Chaining for an empirical process).

### v3

@@ 1. C’7)
    Inverse Continuity Modulus: Let $f_{1},f_{2}\in\mathcal{F}$ for some $\alpha>2+\frac{d}{2}$. Assume there exists a constant $\operatorname{B}\in\mathbb{R}_{>0}$, such that $\|f_{1}\|_{C^{1}(\mathcal{O})}\lor\|G^{D}(f_{1})\|_{C^{2}(\mathcal{O})}\leq\operatorname{B}$ and additionally that the heat source function $g$ satisfies $\inf_{x\in\mathcal{O}}g(x)\geq g_{\min}$. Lemma 24 in Nickl et al. (2020) then provides the existence of a finite constant $c_{1}=c_{1}(g_{\min},f_{\min},\operatorname{B},\mathcal{O},d)\in\mathbb{R}_{>0}$, such that
    $\|f_{1}-f_{2}\|_{\mathbb{L}^{2}(\mathcal{O})}\leq c_{1}\cdot\|f_{2}\|_{C^{1}(\mathcal{O})}\cdot\|G^{D}(f_{1})-G^{D}(f_{2})\|_{H^{2}(\mathcal{O})}.$
    By an application of eq. 27 with $\beta=2$, we further have
    $\|G^{D}(f_{1})-G^{D}(f_{2})\|_{H^{2}(\mathcal{O})}\\ \leq c_{2}(\alpha,\mathcal{O})\|G^{D}(f_{1})-G^{D}(f_{2})\|_{H^{\alpha+1}(\mathcal{O})}^{\frac{2}{\alpha+1}}\cdot\|G^{D}(f_{1})-G^{D}(f_{2})\|_{\mathbb{L}^{2}(\mathcal{O})}^{\frac{\alpha-1}{\alpha+1}}.$
    Moreover, by an application of Lemma 23 in Nickl et al. (2020), we have
    $\|G^{D}(f_{1})-G^{D}(f_{2})\|_{H^{\alpha+1}(\mathcal{O})}^{\frac{2}{\alpha+1}}\leq c_{3}\cdot(1+\|f_{1}\|_{H^{\alpha}(\mathcal{O})}^{2\alpha}\lor\|f_{2}\|_{H^{\alpha}(\mathcal{O})}^{2\alpha})$
    with a finite constant $c_{3}=c_{3}(\alpha,d,\mathcal{O},f_{\min},\|g\|_{H^{\alpha-1}(\mathcal{O})})\in\mathbb{R}_{>0}$. Thus, we have
    $\|f_{1}-f_{2}\|_{\mathbb{L}^{2}(\mathcal{O})}\leq c_{4}\cdot(1+\|f_{1}\|_{H^{\alpha}(\mathcal{O})}^{2\alpha}\lor\|f_{2}\|_{H^{\alpha}(\mathcal{O})}^{2\alpha})\cdot\|G^{D}(f_{1})-G^{D}(f_{2})\|_{\mathbb{L}^{2}(\mathcal{O})}^{\frac{\alpha-1}{\alpha+1}}$
    for a constant $c_{4}=c_{4}(\alpha,d,\mathcal{O},f_{\min},g_{\min},\operatorname{B},\|g\|_{\alpha-1})\in\mathbb{R}_{>0}$.
    Now, let $\operatorname{M}\in\mathbb{R}_{>0}$ be arbitrary and let $f_{1}$ and $f_{2}$ be chosen in such a way that $\|f_{1}\|_{H^{\alpha}(\mathcal{O})}+\|f_{2}\|_{H^{\alpha}(\mathcal{O})}\leq\operatorname{M}$ and $\|G^{D}(f_{1})-G^{D}(f_{2})\|_{\mathbb{L}^{2}(\mathcal{O})}\leq\delta$ for $\delta\in\mathbb{R}_{>0}$ small enough. By an application of Lemma 23 in Nickl et al. (2020) and the embedding $H^{\alpha}(\mathcal{O})\hookrightarrow C^{1}(\mathcal{O})$, we have $\|f_{1}\|_{C^{1}(\mathcal{O})}\lor\|G^{D}(f_{1})\|_{C^{2}(\mathcal{O})}\leq c(\alpha,d,\mathcal{O},f_{\min},\|g\|_{H^{\alpha-1}(\mathcal{O})})$. By above computations we derive
    $\|f_{1}-f_{2}\|_{\mathbb{L}^{2}(\mathcal{O})}\leq c_{5}\cdot\delta^{\frac{\alpha-1}{\alpha+1}}$
    with $c_{5}=c_{5}(\alpha,d,\mathcal{O},f_{\min},g_{\min},\operatorname{M},\|g\|_{\alpha-1})\in\mathbb{R}_{>0}$. Taking the supremum over all admissible $f_{1}$ and $f_{2}$ as in the spirit of 2.9 shows the claim with $\tau=\frac{\alpha-1}{\alpha+1}\in(0,1)$.
We are now providing the proof of theorem 3.2.
Proof of theorem 3.2.
@@
    Thus, for every $\bar{c}\in\mathbb{R}_{>0}$ there exists a constant $c_{1}=c_{1}(\alpha,f_{\min},\|g\|_{\infty},d,\mathcal{O},\sigma)\in\mathbb{R}_{>0}$ sufficiently large, such that
    $\mathbb{P}_{f_{\mathfrak{o}}}^{N}\left(\mathfrak{D}_{\mathfrak{r}}^{2}(\hat{f}_{\textsc{MAP}},f_{\mathfrak{o}})\geq c_{1}(\mathfrak{r}_{N}^{2}\|\Psi^{-1}\circ f_{\mathfrak{o}}\|_{H^{\alpha}(\mathcal{O})}^{2}+m^{2})\right)\leq c_{2}\exp\leq\left(-\bar{c}m^{2}N\right)$
    for all $m\geq\mathfrak{r}_{N}$ satisfying eq. C4 and eq. C5. We identify $f_{\mathfrak{o}}\in\mathcal{F}$ with some $\theta_{\mathfrak{o}}=\Psi^{-1}\circ f_{\mathfrak{o}}\in H^{\alpha}_{c}(\mathcal{O})$ and conclude $\|\Psi^{-1}\circ f_{\mathfrak{o}}\|_{H^{\alpha}(\mathcal{O})}\leq c_{3}(\Psi,\operatorname{L})\in\mathbb{R}_{>0}$ uniformly over all $f_{\mathfrak{o}}\in\mathcal{F}(\operatorname{L})$. Choose $\operatorname{M}\geq 1\lor c_{3}$ and $m:=\operatorname{M}\mathfrak{r}_{N}\geq\mathfrak{r}_{N}$, such that $c_{1}(\mathfrak{r}_{N}^{2}\|\Psi^{-1}\circ f_{\mathfrak{o}}\|_{H^{\alpha}(\mathcal{O})}^{2}+m^{2})\leq 2c_{1}\operatorname{M}^{2}\mathfrak{r}_{N}^{2}$. Then,
    $\mathbb{P}_{f_{\mathfrak{o}}}^{N}\left(\mathfrak{D}_{\mathfrak{r}}^{2}(\hat{f}_{\textsc{MAP}},f_{\mathfrak{o}})\geq 2c_{1}\operatorname{M}^{2}\mathfrak{r}_{N}^{2}\right)\leq c_{2}\exp\leq\left(-\bar{c}M^{2}N\mathfrak{r}_{N}^{2}\right).$
@@
  3. iii)
We are now providing the proof of theorem 3.3.
Proof of theorem 3.3.
The proof follows the lines of the proofs of analogous results in the White Noise Model setting derived in Nickl et al. (2020) (see Proof of Theorem 9). Hence, in this work we only provide a sketch of the proof and ommit the details, which are left for the interested reader.
@@
    By an application of eq. 27, we obtain
    $\|G^{D}(\hat{f}_{\textsc{MAP}})-G^{D}(f_{\mathfrak{o}})\|_{H^{\beta}(\mathcal{O})}^{2}\\ \leq c_{1}\cdot\|G^{D}(\hat{f}_{\textsc{MAP}})-G^{D}(f_{\mathfrak{o}})\|_{\mathbb{L}^{2}(\mathcal{O})}^{\frac{2(\alpha+1-\beta)}{\alpha+1}}\cdot\|G^{D}(\hat{f}_{\textsc{MAP}})-G^{D}(f_{\mathfrak{o}})\|_{H^{\alpha+1}(\mathcal{O})}^{\frac{2\beta}{\alpha+1}}.$
    The first factor can be directly upper bounded by $\mathfrak{D}_{\mathfrak{r}}^{2}(\hat{\theta}_{\textsc{MAP}},\theta_{\mathfrak{o}})$. The second factor can be further upper bounded by firstly applying Lemma 23 in Nickl et al. (2020) and then by considering the definition of $\mathfrak{D}_{\mathfrak{r}}^{2}(\hat{\theta}_{\textsc{MAP}},\theta_{\mathfrak{o}})$. Decomposing the underlying probability space into slices
    $A_{0}:=\left\{\mathfrak{D}_{\mathfrak{r}}^{2}(\hat{f}_{\textsc{MAP}},f_{\mathfrak{o}})<\delta_{N}^{2}\right\}\text{ and }A_{j}:=\left\{\mathfrak{D}_{\mathfrak{r}}^{2}(\hat{f}_{\textsc{MAP}},f_{\mathfrak{o}})\in(2^{2(j-1)}\delta_{N}^{2},2^{2j}\delta_{N}^{2}]\right\}$
    for some $\delta_{N}\propto\mathfrak{r}_{N}$, the corresponding expectation can be upper bounded by
    $\mathbb{E}_{f_{\mathfrak{o}}}^{N}\left[\|G^{D}(\hat{f}_{\textsc{MAP}})-G^{D}(f_{\mathfrak{o}})\|_{H^{\beta}(\mathcal{O})}^{2}\right]\\ \lesssim\sum_{j\in\mathbb{N}_{0}}\mathbb{E}_{f_{\mathfrak{o}}}^{N}\left[\mathds{1}_{A_{j}}\mathfrak{D}_{\mathfrak{r}}^{2}(\hat{\theta}_{\textsc{MAP}},\theta_{\mathfrak{o}})^{\frac{\alpha+1-\beta}{\alpha+1}}\left(1+(\mathfrak{r}_{N}^{-2}\mathfrak{D}_{\mathfrak{r}}^{2}(\hat{\theta}_{\textsc{MAP}},\theta_{\mathfrak{o}}))^{\alpha^{2}\beta}\right)\right]\\ \lesssim\delta_{N}^{\frac{2(\alpha+1-\beta)}{\alpha+1}}\bigg(1+\sum_{j\in\mathbb{N}_{0}}2^{2j\frac{\alpha+1-\beta}{\alpha+1}}(1+2^{2j\alpha^{2}\beta})\mathbb{P}_{f_{\mathfrak{o}}}^{N}(A_{j})\bigg).$
    Using the concentration inequality of theorem 3.2, we obtain
    $\mathbb{E}_{f_{\mathfrak{o}}}^{N}\left[\|G^{D}(\hat{f}_{\textsc{MAP}})-G^{D}(f_{\mathfrak{o}})\|_{H^{\beta}(\mathcal{O})}^{2}\right]\lesssim\delta_{N}^{\frac{2(\alpha+1-\beta)}{\alpha+1}}\cdot(1+o(N^{-\frac{1}{2}})),$
@@
  2. ii)
    Similiarly, as in the proof of C’7) before, we upper bound
    $\|\hat{f}_{\textsc{MAP}}-f_{\mathfrak{o}}\|_{\mathbb{L}^{2}(\mathcal{O})}^{2}\lesssim\|\hat{f}_{\textsc{MAP}}\|_{C^{1}(\mathcal{O})}^{2}\cdot\|G^{D}(\hat{f}_{\textsc{MAP}})-G^{D}(f_{\mathfrak{o}})\|_{H^{2}(\mathcal{O})}^{2}.$
    Thus, by an application of eq. 27 with $\beta=2$, we obtain
    $\|G^{D}(\hat{f}_{\textsc{MAP}}-G^{D}(f_{\mathfrak{o}})\|_{H^{2}(\mathcal{O})}^{2}\\ \lesssim\|G^{D}(\hat{f}_{\textsc{MAP}}-G^{D}(f_{\mathfrak{o}})\|_{H^{1+\alpha}(\mathcal{O})}^{\frac{4}{1+\alpha}}\|G^{D}(\hat{f}_{\textsc{MAP}})-G^{D}(f_{\mathfrak{o}})\|_{\mathbb{L}^{2}(\mathcal{O})}^{\frac{2(\alpha-1)}{\alpha+1}}.$
    Using the same techniques as for the prediction error, we obtain the bound
    $\mathbb{E}_{f_{\mathfrak{o}}}^{N}\left[\|\hat{f}_{\textsc{MAP}}-f_{\mathfrak{o}}\|_{\mathbb{L}^{2}(\mathcal{O})}^{2}\right]\lesssim\delta_{N}^{\frac{2(\alpha-1)}{\alpha+1}}\cdot(1+o(N^{-\frac{1}{2}})),$
@@
4.3 Auxiliary results
The following lemma was proven in Nickl and Wang (2024) (see Lemma 3.12) and is a key tool in the proof of theorem 2.6.
Lemma § 4.7 (Chaining for an Empirical Process).